OpenStudy (anonymous):

The area A enclosed by a square, in square inches, is a function of the length of one of its sides x, when measured in inches. this relation is expressed by the formula A(x)=x^2 for x>0. find A(3) and solve A(x)=36 interpret your answers to each

OpenStudy (jdoe0001):

 if x = 3, what's the Area? if the Area = 36, what's "x"?

OpenStudy (anonymous):

ah ok A(3)=9

OpenStudy (jdoe0001):

right

OpenStudy (anonymous):

flutterer

OpenStudy (jdoe0001):

\(\bf Area = x^2\qquad \qquad 36 = x^2 \qquad \textit{ what's "x"?}\)

OpenStudy (anonymous):

\[\pm6\]

OpenStudy (jdoe0001):

:)